OpenStudy (anonymous):

A worker at a computer factory can assemble 8 computers per hour. How long would it take this worker to assemble 200 computers? A. 8 h B. 20 h C. 25 h D. 200 h

OpenStudy (anonymous):

@AnswerMyQuestions C?

OpenStudy (igreen):

Set up a proportion.

OpenStudy (igreen):

Yes, C is correct.

OpenStudy (anonymous):

@iGreen Thanks xD Just wanted to make sure.

OpenStudy (igreen):

\(\sf \dfrac{8}{1} = \dfrac{200}{x}\) Cross multiply: \(\sf 8 \times x = 200 \times 1\) \(\sf 8x = 200\) Divide 8 to both sides: \(\sf x = 25\) \(\sf \Huge\color{lime}\checkmark\)